# A Match Made in Heaven? AI-driven Matching of Vulnerabilities and Security Unit Tests

### v3

@@ 1. Introduction
Security unit tests trigger vulnerabilities with crafted payloads and use assertions to confirm they are really present (Felderer et al., 2016; Cruzes et al., 2017; Mohammadi et al., 2016). Listing 1 shows a security unit test for a path traversal vulnerability (CWE-22) affecting Apache JSPWiki, which is localized in the getForwardPage() method and is disclosed via CVE-2019-0225. Lines 3–4 show the official patch for this vulnerability. If the test method testNastyDoPost() is executed on the vulnerable version of getForwardPage(), it will fail because the focal method does not return the main wiki page as intended, given the crafted payload at Line 9. Hence, the failed assertion at Line 16 confirms the presence of the vulnerability. A test case behaving like this is called vulnerability-witnessing test (Kang et al., 2022) (a.k.a. Proof of Vulnerability, PoV (Pinconschi et al., 2021; Bui et al., 2022)), or simply “witnessing test” throughout this paper.
@@
Thus, the software security research community is in dire need of more samples of vulnerability witnessing tests. Some compelling use cases for such data are: (i) Understanding the nature of witnessing tests, (ii) training AI models that support the generation of security tests (Kang et al., 2022; Zhang et al., 2023), and (iii) validating the techniques that produce security patches (Bui et al., 2022; Liu et al., 2019; Mohammadi et al., 2019a; Gao et al., 2021). In this paper, we present VuTeCo (VUlnerability TEst COllector), a fully-static AI-driven framework that retrieves vulnerability-witnessing tests in Java test suites. VuTeCo addresses two tasks: (1) The “Finding” task to determine whether a test case is security-related, and (2) the “Matching” task to pair a test case to the specific vulnerability it is witnessing.
VuTeCo has been evaluated in a hold-out set of Vul4J, where it achieved $0.73$ $F_{0.5}$ score and $0.83$ precision in the Finding task. In the Matching task, VuTeCo scored $0.65$ $F_{0.5}$ score and $0.75$ precision. Afterwards, VuTeCo was used in a large-scale mining campaign on GitHub (427 projects), where it retrieved 224 confirmed security-related test cases and 35 tests correctly matched to 29 vulnerabilities. Thus, VuTeCo provides valuable support for large-scale retrieval of security unit test examples for many downstream applications.
In summary, this paper: (1) Introduces VuTeCo, the first ever framework to retrieve vulnerability-witnessing tests in Java repositories; (2) extensively evaluates several AI model types to find the right configuration for VuTeCo; and (3) releases Test4Vul, a dataset containing 259 confirmed security-related unit tests found and matched by VuTeCo in the wild.
@@ 2.3. The Matching Task
The final wording of the system prompt was developed with assistance from ChatGPT (GPT-4o) in May 2025, which helped us refine our initial version (in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025)) to better suit an LLM. The placeholders [System], [User], and [Assistant] are not part of the prompt and only indicate the three roles recognized by DeepSeek Coder. This prompt is further transformed by applying the DeepSeek Coder’s chat template, which starts with the system prompt, then adds ‘### Instruction:’ to initiate the user part, and ‘### Response:’ to start the assistant part. The resulting text is tokenized with Byte-Pair Encoding.
@@ 2.4. Tool-assisted Input Retrieval
Additionally, in the Matching task, the tool also accepts a set of CVE identifiers from the user, which are used to look up an internal knowledge base of 1,992 Java vulnerabilities to retrieve their descriptions. We prepared this catalog by aggregating and de-duplicating three established datasets of Java vulnerabilities, i.e., ProjectKB (SAP, 2025), REEF (Wang et al., 2024b), and ReposVul (Wang et al., 2024a). The catalog also stores additional useful metadata, such as CWE (Common Weakness Enumeration) and known fix commits. If a CVE identifier is not found in the catalog, the tool can retrieve the missing data online using the Vulnerability-Lookup API (Computer Incident Response Center Luxembourg (2025), CIRCL).
Afterward, the tool sends the retrieved data to the model responsible for the requested task (Finding or Matching). In the Finding case, each collected test case is sent to the Finder model to assess its relevance to security. In the Matching case, the test cases are paired with the vulnerability descriptions in all possible ways (i.e., Cartesian product), and each is sent to the Matcher model to assess whether the test case witnesses the paired vulnerability.
@@ 3. Evaluation Design
We conducted a two-phase evaluation. First, we performed an experimental evaluation to search for the best AI model to carry out the Finding and Matching tasks. This involved testing several AI model types on a hold-out set originating from Vul4J—the same source also used for their training. We involved six main models: Three based on pre-trained transformers, CodeBERT (Feng et al., 2020), CodeT5+ (Wang et al., 2023), and UniXcoder (Guo et al., 2022); and three generative large language models, CodeLlama (Rozière et al., 2024), DeepSeek Coder (Guo et al., 2024), and Qwen2.5-Coder (Hui et al., 2024). We hereby distinguish the two groups as code representation models (CRMs) and large language models (LLMs), respectively. The CRMs transform the input into fixed-length embeddings and process them with a linear classification head to obtain the predicted classes. The LLMs, instead, are prompted to analyze the input and provide a binary response directly. We downloaded from Hugging Face the snapshots: codebert-base, codet5p-220m, unixcoder-base, CodeLlama-7b-Instruct-hf, deepseek-coder-6.7b-instruct, Qwen2.5-Coder-7B-Instruct. This evaluation aimed to assess the suitability of such models for the Finding and Matching tasks using validated data. For the Finding task, the six models were trained and tested on a collection of test cases extracted from projects in Vul4J. In the Matching task, the six models were trained and tested using the same set of test cases, but also paired with vulnerability descriptions that affected the corresponding projects. Thus, we carried out a total of 12 distinct train-test sessions.
@@ 3.1.1. Data Selection and Preprocessing
The primary source of data for this experimental evaluation was Vul4J (Bui et al., 2022), as it is the only known source with validated JUnit test methods matched with the vulnerabilities they witness. At the time of this paper writing, Vul4J had 79 reproducible vulnerabilities across 51 Java projects. From these projects, we checked out their patched revisions (i.e., the project versions in which the vulnerability has been fixed and where the witnessing tests have been found) and collected all their test suites using the heuristic described in Section 2.4. We fetched metadata—including descriptions—of the 79 vulnerabilities appearing in Vul4J using the VuTeCo’s internal knowledge base (Section 2.4), resulting in 76 vulnerabilities with metadata (three were reported through inaccessible bug trackers rather than CVE).
For the Finding task, we considered all the 108 JUnit test methods reported Vul4J as confirmed examples of witnessing tests, labeling them as “Security” (the positive class). To create the negative class (“Unclear”), we used all the remaining non-duplicated JUnit test methods mined from the 51 projects in Vul4J, totaling 39,542. We note that the negative class for the Finding task comprises test cases lacking sufficient evidence of their security relevance, rather than being tests that are entirely unrelated to security (hence the name “Unclear”). Then, we split the dataset using stratified sampling (i.e., keeping the same class distribution), creating a training set $TR_{F}$ (70%), a development set $DE_{F}$ (10%), and a test set $TE_{F}$ (20%).
For the Matching task, we created pairs of test cases and the descriptions of the witnessed vulnerabilities in Vul4J. For example, the test method testSendingStringMessage() was paired with the vulnerability CVE-2015-0263 (according to the dataset entry VUL4J-3). Due to the three vulnerabilities discarded previously, the total number of valid pairs was 105, which formed the positive class (“Matched”). To create the negative class (“Not-Matched”), we paired all test methods in Vul4J with unrelated vulnerabilities affecting the same project, forming 84,506 invalid pairs. Just as for the Finding task, we split this dataset with stratified sampling, creating a training set $TR_{M}$ (70%), a development set $DE_{M}$ (10%), and a test set $TE_{M}$ (20%).
@@ 3.1.2. Performance Indicators
VuTeCo’s primary goal is to maximize the number of correct findings while minimizing incorrect ones. The aim is to develop an approach that identifies examples of witnessing tests in software repositories, addressing the challenges outlined in Section 1. Given this circumstance, we include the $F_{0.5}$ score in the analysis, which is a varied version of the $F_{1}$ score where twice as much weight is given to precision compared to recall (Van Rijsbergen, 1979). We prioritized this score in the experimental evaluation because it quantifies the trade-off between precision and recall and aligns with VuTeCo’s requirements. Additional metrics, such as the AUC-ROC, are reported in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025).
@@ 3.1.3. Model Configuration
Hyperparameter Optimization. Since the number of factors for the Finding task was fewer than for the Matching task, we could run additional experiments by optimizing some hyperparameters to further improve performance, selecting the variants with the highest $F_{0.5}$ score and the lowest loss on the development set. We optimized the “intensity” of augmentation mechanisms, i.e., the number of times the JavaTransformer or SPAT were run during oversampling ($5$ and $15$) and the target imbalance ratio for bootstrapping ($0.25$ and $0.50$). The two layers of the CRMs’ classification head were set to $512$ and $128$ in size; however, for the Finding task, we also optimized their size by searching in the space $\{512,256\}$ and $\{128,64\}$, respectively.
@@ 3.1.4. Model Implementation and Training
The experiment was conducted on a server equipped with an NVIDIA Tesla A100 Core GPU and an Intel Xeon Platinum 8352V CPU. Counting all combinations of factors (the details are in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025)), we ran 36 train-testing sessions for the Finding task, of which 24 were for the CRMs and 12 were for the LLMs. On average, each CRM session took 4 hours, whereas each LLM session took 50 hours. For the Matching task, we ran a total of 324 train-testing sessions. For the CRMs, we had 72 sessions from the case without decomposition into two sub-models, while the other 216 sessions were from the case with two sub-models. For the LLMs, we had 36 sessions. On average, each CRM session took 5 hours, whereas each LLM session took 60 hours.
@@ 3.1.5. Baseline Approaches
Baselines for Finding. The approach GrepFind checks if the test method code contains one or more security-specific keywords, such as ‘secur’, ‘cve’, ‘xss’; if so, it is flagged as “Security”. We used the set of keywords defined by Zhou and Sharma (Zhou and Sharma, 2017) and further expanded with additional ones (the full list is in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025)). In essence, this approach behaves like the grep -e command. We conducted experiments varying the required number of matches from $1$ to $5$, testing each condition with and without our extended list, totaling 10 configurations. The VocabFind approach, instead, fits a vocabulary of terms on the test methods in the training set $TR_{F}$. Then, it checks if the terms of the input test method adhere to the fitted vocabulary (i.e., the test method shares a similar vocabulary with the known test cases). If the number of matched terms exceeds the threshold $N$, the input test method is flagged as “Security”. The terms were extracted in two ways: (i) using YAKE, which picks the $K$ most relevant terms from a text in an unsupervised manner (Campos et al., 2020), and (ii) retrieving all the identifiers (i.e., variable and method names) in the test case, as they likely indicate the purpose of the test and are not mixed with other irrelevant language-specific keywords. We refer to these two flavors as VocabFindYAKE and VocabFindIden, respectively. We experimented with NN==$[1..10]$ (step of $1$) for both flavors. For VocabFindYAKE we experimented with KK==$[5..30]$ (step of $5$), while for VocabFindIden we experimented with honoring camelCase and snake_case notations (e.g., whether identifiers like user_Password must be split into user and password), totaling 80 configurations.
@@
Baselines for Matching. The GrepMatch approach checks if the test method code matches one or more terms appearing in the paired vulnerability description; if so, the pair is flagged as “Matched”. The terms are vulnerability descriptions obtained via word-level tokenization and removal of English stop words. We varied the required number of hits from $1$ to $5$. The SimMatch approach checks the similarity between the test method code and the vulnerability description and flags the pair as “Linked” if the similarity score surpasses an arbitrary threshold $T$. We made two different implementations. SimMatchYAKE extracts the keyword sets from both inputs using YAKE (Campos et al., 2020) and compares them using the Jaccard index. SimMatchCRM uses one of three CRMs, i.e., CodeBERT (Feng et al., 2020), CodeT5+ (Wang et al., 2023), and UniXcoder (Microsoft, 2025b), to create embeddings for both inputs and compare them using cosine similarity. For SimMatchYAKE we experimented with TT==$[0.01..0.05]$ (step of $0.01$) and KK==$[5..30]$ (step of $5$). For SimMatchCRM we experimented with five similarity thresholds depending on the embedding models: TT==$[0.91..0.95]$ (step of $0.01$) for CodeBERT, for CodeT5+ TT==$[0.7..0.9]$ (step of $0.05$), and for UniXcoder TT==$[0.3..0.5]$ (step of $0.05$). A total of 45 configurations have been evaluated. Lastly, we developed FixCommits, which inspects the fix commits of a vulnerability $v$, collects the set of added or modified test methods $TM$, and flags the pairs $(t,v),\forall t\in TM$ as “Matched”. Any other pair is considered as “Not-Matched”. This approach relies on the assumption that developers create unit tests alongside the patches to demonstrate that the vulnerability has been successfully fixed. We note that there is currently insufficient evidence to support that this core assumption always holds, as developers may not create tests when fixing vulnerabilities (Bui et al., 2022). For example, in CVE-2010-0684 of ActiveMQ none of the three fix commits (2895197, fed39c3, and 9dc43f3) added any test. Besides, FixCommits can only be run when the fix commit is known, which is not the case for all vulnerabilities disclosed on CVE—while VuTeCo can be used regardless of the fix commit. For these reasons, FixCommits is treated as another heuristic method rather than a ground truth.

### v4

@@ 1. Introduction
Security unit tests trigger vulnerabilities with crafted payloads and use assertions to confirm they are present (Felderer et al., 2016; Cruzes et al., 2017; Mohammadi et al., 2016). Listing 1 shows a unit test for a path traversal vulnerability (CWE-22) affecting Apache JSPWiki, localized in the getForwardPage() method disclosed via CVE-2019-0225. Lines 3–4 show the official patch for this vulnerability. If the test method testNastyDoPost() is run on the vulnerable version of getForwardPage(), it fails because the focal method does not return the main wiki page as intended, due to the crafted payload at Line 9. Hence, the failed assertion at Line 16 confirms the presence of the vulnerability. We refer to any test case that behaves in this way as a vulnerability-witnessing test (Kang et al., 2022) (a.k.a. a Proof of Vulnerability, PoV (Pinconschi et al., 2021; Bui et al., 2022)), or simply ‘witnessing test’ throughout this paper. These tests differ from ordinary test cases in that they focus on identifying security flaws rather than purely functional defects.
@@
Thus, the software security research community is in dire need of more samples of vulnerability witnessing tests. Some compelling use cases for such data are: (i) distilling knowledge about the structure of typical tests for each vulnerability type, (ii) training AI models that can generate or improve security tests (Kang et al., 2022; Zhang et al., 2023), (iii) verifying the correctness of patches produced by Automated Vulnerability Repair tools (Bui et al., 2022; Liu et al., 2019; Mohammadi et al., 2019a; Gao et al., 2021). In this paper, we present VuTeCo (VUlnerability TEst COllector), a fully static AI-driven framework that retrieves vulnerability-witnessing tests from Java test suites. VuTeCo addresses two tasks: (1) the “Finding” task to determine whether a test case is security-related, and (2) the “Matching” task to pair a test case to the specific vulnerability it is witnessing.
VuTeCo has been evaluated in a hold-out set of Vul4J, where it achieved $0.73$ $F_{0.5}$ score and $0.83$ precision in the Finding task. In the Matching task, VuTeCo scored $0.65$ $F_{0.5}$ score and $0.75$ precision. Afterwards, VuTeCo was used in a large-scale mining campaign on GitHub (427 projects), where it retrieved 224 confirmed security-related test cases, and 35 tests correctly matched to 29 vulnerabilities. VuTeCo provides valuable support for large-scale retrieval of security unit test examples for many downstream applications.
In summary, this paper: (1) introduces VuTeCo, the first ever framework to retrieve vulnerability-witnessing tests in Java repositories; (2) extensively evaluates several AI model types to find the right configuration for VuTeCo; and (3) releases Test4Vul, a dataset containing 259 confirmed security-related unit tests found and matched by VuTeCo in the wild. Data and scripts of this study are released in an online replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026) (see Section “Data Availability”).
@@ 2.3. The Matching Task
The final wording of the system prompt was developed with assistance from ChatGPT (GPT-4o) in May 2025, which helped us refine our initial version (in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026)) to better suit an LLM. The placeholders [System], [User], and [Assistant] are not part of the prompt and only indicate the three roles recognized by DeepSeek Coder. This prompt is further transformed by applying the DeepSeek Coder’s chat template, which starts with the system prompt, then adds ‘### Instruction:’ to initiate the user part, and ‘### Response:’ to start the assistant part. The resulting text is tokenized with Byte-Pair Encoding.
@@ 2.4. Tool-assisted Input Retrieval
Additionally, in the Matching task, the user can supply the tool with a list of CVE identifiers, which are used to look up an internal knowledge base of 1,992 historical Java vulnerabilities to fetch their descriptions. This catalog aggregates and de-duplicates three established datasets of Java vulnerabilities, i.e., ProjectKB (SAP, 2025), REEF (Wang et al., 2024b), and ReposVul (Wang et al., 2024a). The catalog also stores additional useful metadata, such as CWE (Common Weakness Enumeration) and known fix commits. If a CVE identifier is not found in the knowledge base, the tool searches for the missing data on the fly via the Vulnerability-Lookup API (Computer Incident Response Center Luxembourg (2025), CIRCL) and updates the catalog locally.
Afterward, the tool sends the retrieved data to the model responsible for the requested task. In the Finding case, each collected test case is sent to the Finder model to assess its relevance to security. In the Matching case, the test cases are paired with the vulnerability descriptions in all possible ways (i.e., the Cartesian product), and each pair is sent to the Matcher model to assess whether the test case witnesses the corresponding vulnerability. Thus, the ideal use of this tool is to list all CVE identifiers that have ever affected the target project, enabling it to assess every possible match.
@@ 3. Evaluation Design
We conducted a two-phase evaluation. First, we performed an experimental evaluation to search for the best AI model to carry out the Finding and Matching tasks. This involved testing several AI model types on a hold-out set originating from Vul4J—the same source also used for their training. We involved six main models: three based on pre-trained transformers, CodeBERT (Feng et al., 2020), CodeT5+ (Wang et al., 2023), and UniXcoder (Guo et al., 2022); and three generative large language models, CodeLlama (Rozière et al., 2024), DeepSeek Coder (Guo et al., 2024), and Qwen2.5-Coder (Hui et al., 2024). We hereby distinguish the two groups as code representation models (CRMs) and large language models (LLMs), respectively. The CRMs transform the input into fixed-length embeddings and process them with a linear classification head to obtain the predicted classes. The LLMs, instead, are prompted to analyze the input and provide a binary response directly. We downloaded from Hugging Face the snapshots: codebert-base, codet5p-220m, unixcoder-base, CodeLlama-7b-Instruct-hf, deepseek-coder-6.7b-instruct, Qwen2.5-Coder-7B-Instruct. This evaluation aimed to assess the suitability of such models for the Finding and Matching tasks using validated data. For the Finding task, the six models were trained and tested on a collection of test cases extracted from projects in Vul4J. In the Matching task, the six models were trained and tested on the same set of test cases, but also paired with vulnerability descriptions that affected the corresponding projects. Thus, we conducted 12 distinct train-test sessions.
@@ 3.1.1. Data Selection and Preprocessing
The primary source of data for this experimental evaluation was Vul4J (Bui et al., 2022), as it is the only known source with validated JUnit test methods matched with the vulnerabilities they witness. At the time of this paper writing, Vul4J had 79 reproducible vulnerabilities across 51 Java projects. From these projects, we checked out their patched revisions (i.e., the project versions in which the vulnerability has been fixed and where the witnessing tests have been found) and collected all their test suites using the heuristic described in Section 2.4. We fetched metadata—including descriptions—of the 79 vulnerabilities appearing in Vul4J using VuTeCo’s internal knowledge base (Section 2.4), resulting in 76 vulnerabilities with metadata (three were reported through inaccessible bug trackers rather than CVE).
For the Finding task, we considered all the 108 JUnit test methods reported Vul4J as confirmed examples of witnessing tests, labeling them as “Security” (the positive class). To create the negative class (“Unclear”), we used all the remaining non-duplicated JUnit test methods mined from the 51 projects in Vul4J (using the heuristic described in Section 2.4), totaling 39,542. We note that the negative class for the Finding task comprises test cases lacking sufficient evidence of their security relevance, rather than being tests that are entirely unrelated to security (hence the name “Unclear”). Then, we split the dataset using stratified sampling (i.e., keeping the same class distribution), creating a training set $TR_{F}$ (70%), a development set $DE_{F}$ (10%), and a test set $TE_{F}$ (20%).
For the Matching task, we created pairs of test cases and the descriptions of the witnessed vulnerabilities in Vul4J. For example, the test method testSendingStringMessage() was paired with the vulnerability CVE-2015-0263 (according to the dataset entry VUL4J-3). Due to the three vulnerabilities discarded previously, the total number of valid pairs was 105, which formed the positive class (“Matched”). To create the negative class (“Not-Matched”), we paired all test methods in the projects in Vul4J with unrelated vulnerabilities from the same project, producing 84,506 invalid pairs. Akin to the Finding task, we split this dataset using stratified sampling, creating a training set $TR_{M}$ (70%), a development set $DE_{M}$ (10%), and a test set $TE_{M}$ (20%).
@@ 3.1.2. Performance Indicators
VuTeCo’s primary goal is to maximize the number of correct findings while minimizing incorrect ones. The aim is to develop an approach that identifies examples of witnessing tests in software repositories, addressing the challenges outlined in Section 1. Given this circumstance, we include the $F_{0.5}$ score in the analysis, which is a varied version of the $F_{1}$ score where twice as much weight is given to precision compared to recall (Van Rijsbergen, 1979). We prioritized this score in the experimental evaluation because it quantifies the trade-off between precision and recall and aligns with VuTeCo’s requirements. Additional metrics, such as the AUC-ROC, are reported in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026).
@@ 3.1.3. Model Configuration
Hyperparameter Optimization. Since the number of factors for the Finding task was fewer than for the Matching task, we could run additional experiments by optimizing some hyperparameters to further improve performance, selecting the variants with the highest $F_{0.5}$ score and the lowest loss on the development set. We optimized the “intensity” of augmentation mechanisms, i.e., the number of times the JavaTransformer or SPAT were run during oversampling ($5$ and $15$) and the target imbalance ratio for bootstrapping ($0.25$ and $0.50$). The size of the two layers of the CRMs’ classification head was set to $512$ and $128$; however, for the Finding task, we also optimized their size by searching in the space $\{512,256\}$ and $\{128,64\}$, respectively.
@@ 3.1.4. Model Implementation and Training
The experiment was conducted on a server equipped with an NVIDIA Tesla A100 Core GPU and an Intel Xeon Platinum 8352V CPU. Counting all combinations of factors (the details are in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026)), we ran 36 train-testing sessions for the Finding task, of which 24 were for the CRMs and 12 were for the LLMs. On average, each CRM session took 4 hours, whereas each LLM session took 50 hours. For the Matching task, we ran a total of 324 train-testing sessions. For the CRMs, we had 72 sessions from the case without decomposition into two sub-models, while the other 216 sessions were from the case with two sub-models. For the LLMs, we had 36 sessions. On average, each CRM session took 5 hours, whereas each LLM session took 60 hours.
@@ 3.1.5. Baseline Approaches
Baselines for Finding. The approach GrepFind checks if the test method code contains one or more security-specific keywords, such as ‘secur’, ‘cve’, ‘xss’; if so, it is flagged as “Security”. We used the set of keywords defined by Zhou and Sharma (Zhou and Sharma, 2017) and further expanded with additional ones (the full list is in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026)). In essence, this approach behaves like the grep -e command. We conducted experiments varying the required number of matches from $1$ to $5$, testing each condition with and without our extended list, totaling 10 configurations. The VocabFind approach, instead, fits a vocabulary of terms on the test methods in the training set $TR_{F}$. Then, it checks if the terms of the input test method adhere to the fitted vocabulary (i.e., the test method shares a similar vocabulary with the known test cases). If the number of matched terms exceeds the threshold $N$, the input test method is flagged as “Security”. The terms were extracted in two ways: (i) using YAKE, which picks the $K$ most relevant terms from a text in an unsupervised manner (Campos et al., 2020), and (ii) retrieving all the identifiers (i.e., variable and method names) in the test case, as they likely indicate the purpose of the test and are not mixed with other irrelevant language-specific keywords. We refer to these two flavors as VocabFindYAKE and VocabFindIden, respectively. We experimented with NN==$[1..10]$ (step of $1$) for both flavors. For VocabFindYAKE we experimented with KK==$[5..30]$ (step of $5$), while for VocabFindIden we experimented with honoring camelCase and snake_case notations (e.g., if identifiers like user_Password must be split into user and password), totaling 80 configurations.
@@
Baselines for Matching. The GrepMatch approach checks if the test method code matches one or more terms appearing in the paired vulnerability description; if so, the pair is flagged as “Matched”. The terms are vulnerability descriptions obtained via word-level tokenization and removal of English stop words. We varied the required number of hits from $1$ to $5$. The SimMatch approach checks the similarity between the test method code and the vulnerability description and flags the pair as “Linked” if the similarity score surpasses an arbitrary threshold $T$. We made two different implementations. SimMatchYAKE extracts the keyword sets from both inputs using YAKE (Campos et al., 2020) and compares them using the Jaccard index. SimMatchCRM uses one of three CRMs, i.e., CodeBERT (Feng et al., 2020), CodeT5+ (Wang et al., 2023), and UniXcoder (Microsoft, 2025b), to create embeddings for both inputs and compare them using cosine similarity. For SimMatchYAKE we experimented with TT==$[0.01..0.05]$ (step of $0.01$) and KK==$[5..30]$ (step of $5$). For SimMatchCRM we experimented with five similarity thresholds depending on the embedding models: TT==$[0.91..0.95]$ (step of $0.01$) for CodeBERT, for CodeT5+ TT==$[0.7..0.9]$ (step of $0.05$), and for UniXcoder TT==$[0.3..0.5]$ (step of $0.05$). A total of 45 configurations have been evaluated. Lastly, we developed FixCommits, which inspects the fix commits of a vulnerability $v$, collects the set of added or modified test methods $TM$, and flags the pairs $(t,v),\forall t\in TM$ as “Matched”. Any other pair is considered as “Not-Matched”. This approach relies on the assumption that developers create unit tests alongside the patches to demonstrate that the vulnerability has been successfully fixed. We note that there is currently insufficient evidence to support that this core assumption always holds, as developers may not create tests when fixing vulnerabilities (Bui et al., 2022). For example, in CVE-2010-0684 of ActiveMQ none of the three fix commits (2895197, fed39c3, and 9dc43f3) added any test. Besides, FixCommits can only be run when the fix commit is known, which is not the case for all vulnerabilities disclosed on CVE, whereas VuTeCo can be used regardless of the fix commit. Hence, FixCommits is treated as a heuristic method rather than a ground truth.